OpenStudy (anonymous):

I need to simplify the following expression. sqrt (-98)sqrt (-50)

OpenStudy (maheshmeghwal9):

\[\color{red}{\sqrt{-98} \times {\sqrt{-50}}=i \sqrt{98\times50.}}\]

OpenStudy (anonymous):

@maheshmeghwal9 not exactly \[\sqrt{-98} \times \sqrt{-50} = i\sqrt{98} \times i\sqrt{50} = i^2 \sqrt{98 \times 50} = -\sqrt{98 \times 50}\]

OpenStudy (maheshmeghwal9):

\[=>i ^2\sqrt{4900}=i^2*70=70i^2=-70\]

OpenStudy (maheshmeghwal9):

sorry there was i^2

OpenStudy (anonymous):

thank you I got -70 as well

OpenStudy (anonymous):

You just made my day thank you

OpenStudy (maheshmeghwal9):

yw:)

OpenStudy (anonymous):

:) Later have a great night

OpenStudy (maheshmeghwal9):

^_^